# Constraining the curvature-induced quantum gravity scales via gamma-ray bursts

### v1

@@ Abstract
We constrain the parameters that govern curvature-induced quantum gravity time-of-flight (TOF) effects. These TOF delays, which occur due to modified dispersion relations of particles in a vacuum, could be a phenomenological signature of quantum gravity. Gamma-ray bursts (GRBs), short, high-energy events from distant galaxies, offer a unique opportunity to impose observational limits on TOF delays and, by extension, on the energy scales of quantum gravity. Using the standard Jacob-Piran relation, which assumes a locally-flat spacetime, the analysis of quantum gravity-induced TOF effects establishes a lower limit of approximately $10E_{\text{Pl}}$ on the energy scale of these effects. However, curvature-induced quantum gravity effects may introduce additional contributions. From current GRB observations, we find that, at a 95% credibility level, in the symmetry-deformed scenario, curvature-induced TOF effects may only arise at energies above $0.04E_{\rm Pl}$. If we consider only curvature-induced effects, this limit is an order of magnitude stronger. Observing more GRBs at different redshifts could improve the constraints on the curvature-induced QG phenomena. However, given the capabilities of current telescopes and the current understanding of GRBs, it is unlikely that these constraints will be significantly extended beyond the present level.
@@ 1 Introduction
One of the potential phenomenological manifestations of quantum gravity (QG) is a modification of the in-vacuum dispersion relation, i.e., breaking or deformation of the Lorentz invariance of the speed of massless particles [1]. Microscopically, it corresponds to the treatment of the quantum vacuum of the gravitational field as a dispersing medium, which can be polarized by a by-passing particle, with the corresponding back-reaction. The modified dispersion relation (MDR) leads to the time-of-flight (TOF) delay of photons with different energies. It is a cornerstone of astrophysical tests for QG phenomenology [2]: roughly speaking, detecting two photons of different energies from the same event; one can consider the duration of this event as an upper limit for the QG TOF delay between these photons. This allows us to set a lower limit on the energy scale $E_{\text{QG}}$ at which the QG effects on the dispersion relation become significant.
Gamma-ray bursts (GRBs) are the best transients for such purposes. These are short-duration high-energy transients, that are variable on short time scales. They are cosmologically distant, so the QG TOF delay would accumulate up to relatively large values. Thus, QG TOF delay, if exists, would be relatively significant in GRB observations. While there is no explicit detection of QG effects yet, current studies of TOF delays in individual GRBs allow to set limits on the quantum-gravity energy scale $E_{\text{QG}}$ up to ten times the Planck energy $E_{\text{Pl}}=1.22\times 10^{19}\,$eV [3, 4, 5, 6]. The application of the TOF method to GRBs suffers from systematic uncertainties, mainly due to unknown intrinsic lags between photons of different energies, which are difficult to take into account (e.g., [7, 8, 9]). However, we can overcome those by combining observations from different GRBs at different cosmological distances and from different emission regions with GRBs.
@@
In this work, we compare this finite-parametric form to existing GRB observations and constrain these parameters. In section 2, we list the most stringent limits on $E_{\text{QG}}$, obtained from individual GRBs via the JP08 formula for $\Delta t(\Delta E,z)$. In section 3, we describe the formalism for the QG TOF delay developed by Amelino-Camelia et al. (2024; hereafter A24) [14] that accounts for curvature-induced QG effects. We show that the limits on $E_{\text{QG}}$ obtained from the standard analysis can be transformed to the limits on the parameters of the A24 formula without reprocessing the observational data. In section 4, we set rigorous constraints on these parameters using the Bayesian methodology. We explore, in section 5, which future GRB observations may improve the present results. In the concluding section 6, we briefly summarize our results and describe potential caveats of the analysis performed.
@@ 2 Standard gamma-ray burst tests for QG phenomenology
3 Curvature-induced effects within symmetry-deformed scenario
@@ 3.1 Formalism
The method proposed by A24 [14] to study the curvature-induced effects on the TOF delay consists of two steps. First, one has to develop the formalism of the deformed-symmetry scenario in the de Sitter universe, the limiting case of expanding spacetime with the Hubble constant being exactly constant. The most general deformation of de Sitter algebra of symmetries is five-parametric, but only three independent combinations of them control $\Delta t$. The second step, denoted de Sitter slicing [12], is to generalize the results for the de Sitter universe to a universe with a time-dependent Hubble constant, namely the Friedman-Lemaître-Robertson-Walker (FLRW) universe.
@@ 3.2 Simple estimates of $\eta_{\text{lf}}$, $\eta_{\text{ci1}}$, and $\eta_{\text{ci2}}$
Figure 2: Error bars indicating $95\%$ limits on $\kappa$ from the GRB data for the afterglow (orange) and for the prompt emission (green). The black lines show two limiting examples satisfying all these constraints within the A24 approach. Solid for $\eta_{\text{lf}}=0.52$ and dashed for $\eta_{\text{lf}}=-1.9$, while $\eta_{\text{ci1}}=14.0\eta_{\text{lf}}$ and $\eta_{\text{ci2}}=-11.7\eta_{\text{lf}}$ in both cases. The two right panels are the same as the left panel but are log-log-scaled.
@@ 4 Bayesian constraining of QG parameters
where $\mathcal{N}(x;\mu,\sigma^{2})$ is the normal distribution with mean value $\mu$ and dispersion $\sigma$, the function $\kappa(z,\eta_{\text{lf}},\eta_{\text{ci1}},\eta_{\text{ci2}})$ is defined in eq. (3.4), and $\kappa_{i\,\text{min}}$ and $\kappa_{i\,\text{max}}$ are the lower and upper limits on $\kappa$ for the $i$’th GRB given in table 1. Note that these likelihood profiles approximate the ones provided in the original papers [3, 4, 6]. This may, of course, affect the constraints we obtain. However, comparing the results to the analysis presented earlier suggests that those effects are not significant.
@@
In case there are no curvature-induced TOF effects, that is for the purely locally-flat QG case $\eta_{\text{ci1}}=\eta_{\text{ci2}}=0$, i.e., the JP08 approach, our fitting procedure yields 95% limits on the locally-flat parameter $-0.079<\eta_{\text{lf}}<0.10$. The corresponding constraint on the energy scale is $E_{\text{QG}}^{\text{(lf)}}\gtrsim 10E_{\text{Pl}}$. It is in good agreement with the strongest constraints obtained in previous works [3, 5, 6]. It confirms that the results obtained from different GRBs at different redshifts and from different emission mechanisms (prompt in 090510 and afterglow in 221009A) are consistent with each other.
@@ 6 Conclusion
This work, for the first time, provides a detailed investigation of the constraints on the linear locally-flat and curvature-induced (within the symmetry-deformed scenario [14]) QG TOF effects. Using the A24 formalism, we explored the parameter space for both locally-flat and curvature-induced contributions, setting 95% credibility bounds on all three terms of the generic TOF delay formula. Using observations of the six most suitable GRBs and Bayesian analysis, we find that curvature-induced TOF effects may only emerge at energy scales exceeding $0.04E_{\text{Pl}}$. Because of the additional degrees of freedom introduced, when curvature-induced effects exit, the limit on the locally-flat QG scale is $0.5E_{\text{Pl}}$, as compared to $\sim 10E_{\text{Pl}}$ when these effects don’t exist.
@@
Our findings suggest that future observations can improve these constraints. We identify specific redshift-energy combinations that would maximize the sensitivity to QG parameters. For instance, observations of GeV GRBs, like GRB 090510, at moderate redshifts or nearby TeV GRBs with minimal extragalactic background light absorption could provide valuable insights. However, it is unlikely that the limit of $10E_{\text{Pl}}$ for the locally-flat scale can be improved significantly, as those limits are stretching the capability of both GeV and TeV range telescopes.

### v2

@@ Abstract
We constrain the parameters that govern curvature-induced quantum gravity time-of-flight (TOF) effects. These TOF delays, which occur due to modified dispersion relations of particles in the vacuum, could be a phenomenological signature of quantum gravity. Gamma-ray bursts (GRBs), short, high-energy events from distant galaxies, offer a unique opportunity to impose observational limits on TOF delays and, by extension, on the energy scales of quantum gravity. Using the standard Jacob-Piran relation, which assumes a locally-flat spacetime, the analysis of quantum gravity-induced TOF effects establishes a lower limit of approximately $10E_{\text{Pl}}$ on the energy scale of these effects. However, curvature-induced quantum gravity effects may introduce additional contributions. From current GRB observations, we find that, at a 95% credibility level, in the symmetry-deformed scenario, curvature-induced TOF effects may only arise at energies above $0.04E_{\text{Pl}}$. If we consider only curvature-induced effects, this limit is an order of magnitude stronger. Observing more GRBs at different redshifts could improve the constraints on the curvature-induced QG phenomena. However, given the capabilities of current telescopes and the current understanding of GRBs, it is unlikely that these constraints will be significantly extended beyond the present level.
@@ 1 Introduction
One of the potential phenomenological manifestations of quantum gravity (QG) is a modification of the in-vacuum dispersion relation, i.e., breaking or deformation of the Lorentz invariance of the speed of massless particles [1]. Microscopically, it corresponds to the treatment of the quantum vacuum of the gravitational field as a dispersing medium, which can be polarized by a by-passing particle, with the corresponding back-reaction. The modified dispersion relation (MDR) leads to the time-of-flight (TOF) delay of photons with different energies. It is a cornerstone of astrophysical tests for QG phenomenology [2]: roughly speaking, detecting two photons of different energies from the same event, one can consider the duration of this event as an upper limit for the QG TOF delay between these photons. This allows us to set a lower limit on the energy scale $E_{\text{QG}}$ at which the QG effects on the dispersion relation become significant.
Gamma-ray bursts (GRBs) are the best transients for such purposes. They are short-duration high-energy transients, that are variable on short time scales. They are cosmologically distant, so the QG TOF delay would accumulate up to relatively large values. Thus, QG TOF delay, if exists, would be relatively significant in GRB observations. While there is no explicit detection of QG effects yet, current studies of TOF delays in individual GRBs allow to set limits on the quantum-gravity energy scale $E_{\text{QG}}$ up to ten times the Planck energy $E_{\text{Pl}}=1.22\times 10^{19}\,$eV [3, 4, 5, 6]. The application of the TOF method to GRBs suffers from systematic uncertainties, mainly due to unknown intrinsic lags between photons of different energies, which are difficult to take into account (e.g., [7, 8, 9]). However, we can overcome those by combining observations from different GRBs at different cosmological distances and from different emission regions with GRBs.
@@
In this work, we compare this finite-parametric form to existing GRB observations and constrain these parameters. In section 2, we list the most stringent limits on $E_{\text{QG}}$, obtained from individual GRBs via the JP08 formula for $\Delta t(\Delta E,z)$. In section 3, we describe the formalism for the QG TOF delay developed by Amelino-Camelia et al. (2024; hereafter A24) [14] that accounts for curvature-induced QG effects. We show that the limits on $E_{\text{QG}}$ obtained from the standard analysis can be transformed to the limits on the parameters of the A24 formula without reprocessing the observational data. In section 4, we set rigorous constraints on these parameters using the Bayesian methodology. Finally, we explore, in section 5, which future GRB observations may improve the present results. In the concluding section 6, we briefly summarize our results and describe potential caveats of the analysis performed.
@@ 2 Standard gamma-ray burst tests for QG phenomenology
3 Curvature-induced effects within the symmetry-deformed scenario
@@ 3.1 Formalism
The method proposed by A24 [14] to study curvature-induced effects on the TOF delay consists of two steps. First, one has to develop the formalism of the deformed-symmetry scenario in the de Sitter universe, the limiting case of expanding spacetime with the Hubble constant being exactly constant. The most general deformation of de Sitter algebra of symmetries is five-parametric, but only three independent combinations of them control $\Delta t$. The second step, denoted de Sitter slicing [12], is to generalize the results for the de Sitter universe to a universe with a time-dependent Hubble constant, namely the Friedman-Lemaître-Robertson-Walker (FLRW) universe.
@@ 3.2 Simple estimates of $\eta_{\text{lf}}$, $\eta_{\text{ci1}}$, and $\eta_{\text{ci2}}$
Figure 2: Error bars indicating $95\%$ limits on $\kappa$ from the GRB data for the afterglow (orange) and for the prompt emission (green). The thick black lines show two limiting examples satisfying all these constraints within the A24 approach, solid for $\eta_{\text{lf}}=0.52$ and dashed for $\eta_{\text{lf}}=-1.9$, while $\eta_{\text{ci1}}=14.0\eta_{\text{lf}}$ and $\eta_{\text{ci2}}=-11.7\eta_{\text{lf}}$ in both cases. The thin gray lines show the limiting examples for the case of vanishing locally-flat term, $\eta_{\text{lf}}=0$, $\eta_{\text{ci1}}=\pm 2.5$, and $\eta_{\text{ci2}}=\mp 1.8$. The two right panels are the same as the left panel but are log-log-scaled.
@@
This method can be applied also to the case when $\eta_{\text{lf}}=0$, that is when the locally-flat term vanishes. In this case, $\eta_{\text{ci1}}\propto\eta_{\text{ci2}}$. The constant of proportionality depends on whether we chose GRB 090510 or 221009A as the point in which we set $\kappa\approx 0$. The limits obtained on the curvature-induced terms, that slightly depend on the choice of the anchor point, are $-2.5\lesssim\eta_{\text{ci1}}\lesssim 2.5$ and $-1.8\lesssim\eta_{\text{ci2}}\lesssim 1.8$. The corresponding limiting cases are shown in figure 2 by thin grey lines. Clearly, in view of the strong limits given by the two GRBs 090510 or 221009A the higher redshift points are not significant in this case.
@@ 4 Bayesian constraining of QG parameters
where $\mathcal{N}(x;\mu,\sigma^{2})$ is the normal distribution with mean value $\mu$ and dispersion $\sigma$, the function $\kappa(z,\eta_{\text{lf}},\eta_{\text{ci1}},\eta_{\text{ci2}})$ is defined in eq. (3.4), and $\kappa_{i\,\text{min}}$ and $\kappa_{i\,\text{max}}$ are the lower and upper limits on $\kappa$ for the $i$’th GRB given in table 1. Note that these likelihood profiles approximate the ones provided in the original papers [3, 4, 6]. Still, this approximation preserves the asymmetry, relative to $\kappa=0$, in the data by taking the mean of the Gaussian distribution to be $\mu_{i}=({\kappa_{i\,\text{min}}+\kappa_{i\,\text{max}}})/{2}$, that is half way between the upper and lower limits. This approximation may, of course, affect the constraints we obtain. However, comparing the results (table 2) to those presented in section 3.2, where the upper limits on the QG parameters are dictated solely by the upper error bar of GRB 090926A and the lower ones come from the lower error bar of GRB 090902B, suggests that those effects are not significant.
@@
In case there are no curvature-induced TOF effects, that is for the purely locally-flat QG case $\eta_{\text{ci1}}=\eta_{\text{ci2}}=0$, i.e., the JP08 approach, our fitting procedure yields 95% limits on the locally-flat parameter $-0.079<\eta_{\text{lf}}<0.10$. The corresponding constraint on the energy scale is $E_{\text{QG}}^{\text{(lf)}}\gtrsim 10E_{\text{Pl}}$. It is in good agreement with the strongest constraints obtained in previous works [3, 5, 6]. It confirms that the results obtained from different GRBs at different redshifts and from different emission mechanisms (prompt in 090510 and afterglow in 221009A) are consistent with each other. [^6] One should not be confused by the fact that the $\kappa$ limits for GRB 221009A are much smaller than those for all other bursts. According to Eq. (3.3), the importance of each GRB for estimating the QG scale is defined by $f_{\text{lf}}(z)/\kappa$. Table 1 and figure 1 show that in the locally-flat case the overall constraint is dominated equally by GRB 090510 and GRB 221009A.
@@ 6 Conclusion
This work, for the first time, provides a detailed investigation of the constraints on the linear locally-flat and curvature-induced (within the symmetry-deformed scenario [14]) QG TOF effects. Using the A24 formalism, we explored the parameter space for both locally-flat and curvature-induced contributions, setting 95% credibility bounds on all three terms of the generic TOF delay formula. Using observations of the six most suitable GRBs and Bayesian analysis, we find that curvature-induced TOF effects may only emerge at energy scales exceeding $0.04E_{\text{Pl}}$. Because of the additional degrees of freedom introduced, when curvature-induced effects exist, the limit on the locally-flat QG scale is $0.5E_{\text{Pl}}$, as compared to $\sim 10E_{\text{Pl}}$ when these effects don’t exist.
@@
Our findings suggest that future observations can improve these constraints. Particularly, increasing the number of GRBs suitable for the analysis will make the QG TOF limits more robust with respect to the intrinsic effects problem. We identify specific redshift-energy combinations that would maximize the sensitivity to QG parameters. For instance, observations of GeV GRBs, like GRB 090510, at moderate redshifts, or nearby TeV GRBs with minimal extragalactic background light absorption could provide valuable insights. However, it is unlikely that the limit of $10E_{\text{Pl}}$ for the locally-flat scale can be improved significantly, as those limits are stretching the capability of both GeV and TeV range telescopes.